OpenStudy (anonymous):

Select the rate that is the greatest. A. $100 per 8 hours B. $700 per 50 hours C. $15 per hour D. $25 per 2 hours

OpenStudy (anonymous):

100/8 = 12.5 700/50 = 14 15/1 = 15 25/2 = 12.5 $15 per hour is the greatest :D